# InteractPro: A Unified Framework for Motion-Aware Image Composition

### v2

@@ C.2 Object Placement Decision For InteractMotion
Appendix D More Implementation Details
The InteractPlan module completes in 10 seconds. The image generated by InteractMotion using SVD is rendered at a resolution of 576 1024, whereas the output from the text-to-video-ms-1.7b model (InteractPhys) is generated at 1080 1920 resolution. During InteractMotion, all intermediate composite images from the planning phase and their corresponding masks are reshaped to 576 1024. Instead of using precise object masks, a general bounding box is employed to allow flexibility for object movement in motion-aware composition. InteractMotion takes approximately 90 seconds and consumes 10 GB of memory on a single Tesla V100 GPU with sampling steps. To speed up generation, the number of inference steps can be adjusted, as discussed in Section E.
InteractPhys, on the other hand, takes 60 seconds and uses 15 GB of memory on a single Tesla V100 to render 80 frames, without parameter optimization. With optimization enabled, memory usage increases to 17 GB, and training takes about 40 seconds per epoch, followed by 60 seconds for final rendering. Typically, the system is optimized with 10 epochs.
Appendix E Extended Experiments for InteractMotion
Static Scenarios. While designed for motion-aware compositions, our framework adapts to static scenes by using video diffusion models’ scene understanding and mask restrictions in InteractMotion, preventing hallucination when motion is not expected. The motion parameters in SVD can also be adjusted to limit motion, maintaining identity preservation and seamless blending. Figure 10 demonstrates our method’s versatility in static scene types. In row (a), the bench remains unaffected by the wind, as it is too heavy for interaction with the outdoor breeze. In row (b), the cup remains undisturbed by the indoor environment, as expected.
Speed up generation. The default configuration uses T=25 sampling steps. However, for faster results, the InteractMotion pipeline can be accelerated by reducing T value. As shown in Figure 11, setting T to 15 yields comparable results, cutting generation time from 90 seconds to just 50 seconds without significantly compromising quality.
Appendix F Extra Qualitative Results for InteractPhys
Note that we are unable to compare against the atypical interaction shown in Fig.4 of main paper, as some baselines cannot incorporate additional conditioning—such as text prompts—and can only produce default or typical outputs.
We provide explanation of our physics aware composition InteractPhys in Fig.1(d-f) of main paper. Our method showcases realistic interactions between inserted foreground object and object in given background. In 1(d), the runny egg smoothly spreads over the contours of the hotdog following the shape of hotdog. In 1(e), the branch and leaves of the plant bend under the impact of an apple falling into the plant, while the heavier pot remains stationary. In 1(f), the sandcastle is compressed due to the weight of the book. We also show comparisons with other baselines for these cases in Fig.13. Other methods are unable to demonstrate the smooth flow of runny egg against the hotdog, or changes the appearance of input egg. For the apple example, others only show fruit hanging on the plant and seems like on different image layer or looking like cut and paste, whereas ours perfectly place the apple in between the branches of the plant, causing the branches to tilt. Other methods are also unable to demonstrate the weight impact of the book, which should cause the sandcastle to be compressed.
@@ Appendix F Extra Qualitative Results for InteractPhys
Appendix G Extra Qualitative Results for InteractMotion
We provide explanation of our motion aware composition InteractMotion in Fig.1(a-c) of main paper, with qualitative comparison in Fig.12. Our method showcases realistic interactions between inserted foreground object and object in given background. In (a), a towel is placed on a rope at the beach, where the image captures the towel swaying in the breeze. Our method inherently incorporates scale and rotation transformations, guiding the towel’s natural motion in response to the wind. Similarly in (b), the flag is inserted into the open air grass field, where wind presence is likely, shows the flag swaying naturally in the air. In (c), placing a glass bottle into the ocean gives seamless integration. InteractMotion effectively positions the bottle within the sea layer, avoiding the disjointed appearance typical of Cut and Paste. The realistic distortion from light refraction in water further enhances this effect.

### v3

@@ VII-B Object Placement Decision For InteractMotion
VIII More Implementation Details
The InteractPlan module completes in 10 seconds. The image generated by InteractMotion using SVD is rendered at a resolution of 576 1024, whereas the output from the text-to-video-ms-1.7b model (InteractPhys) is generated at 1080 1920 resolution. During InteractMotion, all intermediate composite images from the planning phase and their corresponding masks are reshaped to 576 1024. Instead of using precise object masks, a general bounding box is employed to allow flexibility for object movement in motion-aware composition. InteractMotion takes approximately 90 seconds and consumes 10 GB of memory on a single Tesla V100 GPU with sampling steps. InteractPhys, on the other hand, takes 60 seconds and uses 15 GB of memory on a single Tesla V100 to render 80 frames, without parameter optimization. With optimization enabled, memory usage increases to 17 GB, and training takes about 40 seconds per epoch, followed by 60 seconds for final rendering. Typically, the system is optimized with 10 epochs.
IX Extra Qualitative Results for InteractPhys
@@
X Extra Qualitative Results for InteractMotion